### v1

@@ Abstract.
Quantum Neural Network (QNN) combines the Deep Learning (DL) principle with the fundamental theory of quantum mechanics to achieve machine learning tasks with quantum acceleration. Recently, QNN systems have been found to manifest robustness issues similar to classical DL systems. There is an urgent need for ways to test their correctness and security. However, QNN systems differ significantly from traditional quantum software and classical DL systems, posing critical challenges for QNN testing. These challenges include the inapplicability of traditional quantum software testing methods, the dependence of quantum test sample generation on perturbation operators, and the absence of effective information in quantum neurons. In this paper, we propose QuanTest, a quantum entanglement-guided adversarial testing framework to uncover potential erroneous behaviors in QNN systems. We design a quantum entanglement adequacy criterion to quantify the entanglement acquired by the input quantum states from the QNN system, along with two similarity metrics to measure the proximity of generated quantum adversarial examples to the original inputs. Subsequently, QuanTest formulates the problem of generating test inputs that maximize the quantum entanglement sufficiency and capture incorrect behaviors of the QNN system as a joint optimization problem and solves it in a gradient-based manner to generate quantum adversarial examples. Experimental results demonstrate that QuanTest possesses the capability to capture erroneous behaviors in QNN systems (generating 67.48%-96.05% more test samples than the random noise under the same perturbation size constraints). The entanglement-guided approach proves effective in adversarial testing, generating more adversarial examples (maximum increase reached 21.32%).
@@ 1. Introduction
Figure 1. The differences between classical DNN testing (upper part) and QNN system testing (lower part). In QNN system testing, the QNN components cannot extract effective neuron information as in classical DNN, and the generation of adversarial examples is also different from the classical approach.
The differences between classical DNN testing (upper part) and QNN system testing (lower part). In QNN system testing, the QNN components cannot extract effective neuron information as in classical DNN, and the generation of adversarial examples is also different from the classical approach.
Figure 1. The differences between classical DNN testing (upper part) and QNN system testing (lower part). In QNN system testing, the QNN components cannot extract effective neuron information as in classical DNN, and the generation of adversarial examples is also different from the classical approach.
@@
    Limited useful information can be extracted from quantum neurons. Owing to the closed and unitary characteristics of quantum systems, apart from the parameters of rotation gates, an individual quantum neuron cannot extract more information. Valuable information often concentrates on the overall output of QNN and the corresponding measurement results. This renders existing DL testing metrics ineffective for QNN systems. Consequently, a novel evaluation metric tailored to quantum systems is imperative for QNN system testing.
In this paper, we propose QuanTest, an adversarial testing framework for QNN systems guided by quantum entanglement. It aims to generate adversarial examples that can trigger more solution spaces in QNN, thereby uncovering more erroneous behaviors in the QNN system. In particular, We first extract information from the quantum system constituted by the QNN model and quantum states. Considering the wave function collapse caused by observing the quantum systems (Nielsen and Chuang, 2010), in addition to measuring the output quantum states from the QNN model to obtain predictive information, it is necessary to retain some copies for subsequent testing to calculate quantitative indicators. For the quantitative metrics, we designed an entanglement adequacy criterion to quantify the entanglement acquired by input quantum states from the QNN system and two similarity metrics to measure the proximity between the generated adversarial examples and the original inputs. Subsequently, we formulate the problem of generating test inputs that maximize quantum entanglement adequacy (QEA) of a QNN system while also exposing as many erroneous behaviors as possible as a joint optimization problem. We effectively solve this problem by obtaining perturbations in a gradient-based manner. The perturbations act directly on the original inputs in the form of a unitary operator, generating quantum adversarial examples through quantum evolution. Finally, we detect quantum adversarial examples and evaluate their quality through model performance and similarity metrics.
@@
  • We introduce quantum entanglement adequacy as the first test adequacy metric for QNN systems, and two similarity metrics to quantify quantum state sample differences.
@@ 2.1.1. Qubits
where $\alpha$ and $\beta$ are complex numbers, representing probability amplitudes and satisfying the condition $|\alpha|^{2}+|\alpha|^{2}=1$. $\{|0\rangle,|1\rangle\}$ consists of two orthogonal 2-dimensional state vectors and is known as the computational basis in this Hilbert space. A composite system of $n$ qubits can be described by a unit complex vector in a $2^{n}$-dimensional Hilbert space. Note that the quantum state space of the composite system is the tensor product of its component system’s state spaces.
@@ 2.1.2. Quantum Logic Gates
where $U^{\dagger}$ is the conjugate transpose of $U$, and $I$ representing the identity matrix. This implies that all quantum gates must be reversible, and the dimensions of input and output must remain consistent to preserve the unitary transformation. A quantum gate acting on $n$ qubits $|\psi\rangle$ can be described by a $2^{n}\times 2^{n}$ unitary matrix $U$. The essence of this operation is to perform matrix operations on the state vector, i.e., $U|\psi\rangle$. Common quantum logic gates can be categorized based on their properties and functions into fixed gates, rotation gates, and controlled gates, among others.
@@ 2.2. Quantum Neural Network
Quantum neural networks (QNNs) are a type of neural network model based on quantum computing, with the core being parameterized quantum circuits (PQCs) (Benedetti et al., 2019) driven by variational quantum algorithms (Cerezo et al., 2021). Similar to classical feedforward neural networks, QNNs learn by updating the trainable parameters of rotation gates within the circuits through gradients——this aspect is executed by classical optimizers, and the processing and computation of the target quantum state are carried out by quantum computers. An example of using a QNN model $\mathcal{M}(|\boldsymbol{x}\rangle;U_{\boldsymbol{\Theta}})$ for machine learning tasks is shown in Figure 2. The training data $\boldsymbol{x}$ is first encoded as quantum states $|\boldsymbol{x}\rangle$. Subsequently, these states are input into the QNN for computation. The trainable parameters $\boldsymbol{\Theta}$ within the PQC $U_{\boldsymbol{\Theta}}$ are updated using classical optimizers until the results meet the expected termination criteria. In this process, the PQC $U_{\boldsymbol{\Theta}}$ can be regarded as a unitary operator, and its action on the quantum state $|\boldsymbol{x}\rangle$ can be represented as $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$. The final output of the QNN can only be obtained after measuring $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$.
@@ 3.1. Quantum Entanglement Adequacy Metrics
One core reason quantum computation can demonstrate potential quantum advantage in many problems is that it can construct an exponentially large quantum state space through controllable entanglement and interference (Havlíček et al., 2019). By utilizing low-depth circuits to generate quantum states with high entanglement characteristics, QNN models may produce atypical patterns that classical models cannot effectively produce. These patterns can efficiently represent solution space for tasks such as prediction and classification, capturing nontrivial correlations in the data (Schuld et al., 2020). Quantum machine learning models such as quantum circuit learning (QCL) (Mitarai et al., 2018), parameterized Hamiltonian learning (PHL) (Shi et al., 2022), and quantum kernel methods (Peters et al., 2021) achieve strong entanglement in quantum circuits for enhancing the entanglement capability of QNNs by incorporating various configurations of multi-qubit gates such as CNOT, iSWAP, and their extended variants. We can perceive that, as the quantum state used as a test input undergoes computation within a QNN, the more adequately it activates entangled neurons, the higher the entanglement degree it obtained from the QNN. This leads to a more comprehensive coverage of the solution space in the QNN. Therefore, we propose a method to calculate the QEA of QNNs. This is achieved by measuring the entanglement of both the test input quantum state and the resulting output quantum state after undergoing unitary operations within the QNN. The difference in measurement outcomes is used to represent the QEA of that particular test case.
@@
Given a fully trained QNN model $\mathcal{M}$ and a test input set $\mathcal{T}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$, the of $\mathcal{M}$ on the input quantum state $|\boldsymbol{x}\rangle$ can be equivalently represented as the multiplication of $|\boldsymbol{x}\rangle$ and a unitary matrix $U_{\boldsymbol{\Theta}}$, where $|\boldsymbol{x}\rangle$ represents the quantum state that encodes the information of $\boldsymbol{x}\in\mathcal{T}$, and $\boldsymbol{\Theta}$ represents the learned and fixed gate parameters in $\mathcal{M}$. Note that the result output of $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$ is a quantum state with the same dimension as $|\boldsymbol{x}\rangle$. Then, the quantum entanglement adequacy (QEA) can be expressed as:
@@
Figure 4. The example of QEA computation for two quantum states in the QNN. The nodes in gray denote quantum entanglement layers that are not sufficiently activated, while aqua nodes represent those that are adequately activated.
The example of QEA computation for two quantum states in the QNN. The nodes in gray denote quantum entanglement layers that are not sufficiently activated, while aqua nodes represent those that are adequately activated.
Figure 4. The example of QEA computation for two quantum states in the QNN. The nodes in gray denote quantum entanglement layers that are not sufficiently activated, while aqua nodes represent those that are adequately activated.
@@
10:         obj_join = [];
11:         obj_joint.append(obj_incorrect($|\boldsymbol{x}\rangle_{out}$, strategy));
12:         obj_joint.append(w $\cdot$ obj_entanglement($|\boldsymbol{x}\rangle_{adv}$, $|\boldsymbol{x}\rangle_{out}$, k));
@@
3.2. QuanTest Algorithm
3.2.1. Joint Optimization
It consists of two parts, where the first part $k\cdot Q(U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle)-Q(|\boldsymbol{x}\rangle)$ is to generate test inputs that maximize QEA, and the second part $f(\mathcal{M}(|\boldsymbol{x}\rangle;U_{\boldsymbol{\Theta}});\mathcal{A})$ is to generate test inputs that trigger as many incorrect behaviors in the QNN as possible. The hyperparameter $\lambda$ is used to balance the two optimization objectives during the joint optimization process. We treat these two parts as a joint optimization problem and maximize this objective function (Algorithm 1 line 8-11).
@@ 3.2.1. Joint Optimization
  • Maximizing incorrect behaviors. Modular construction enables QuanTest to conveniently select different quantum-adapted adversarial attack strategies $\mathcal{A}$. The determination of this strategy dictates the prediction information to be further calculated after measuring the output quantum state during the information extraction. For instance, the quantum-adapted FGSM strategy requires providing loss, while the quantum-adapted DLFuzz strategy needs category probabilities as input. QuanTest aims in this section to modify the input in the form of adversarial attacks, leading the model to exhibit incorrect behavior.
All terms of the jointly optimized objective are differentiable. QuanTest can iteratively modify the input $|\boldsymbol{x}\rangle$ by using gradient ascent to maximize these terms. The computed gradient $\frac{\partial Obj}{\partial|\boldsymbol{x}\rangle}$ will become the perturbation, which is subsequently transformed into a perturbation operator on a quantum circuit for generating adversarial examples.
@@ 3.2.3. Quality Evaluation
The generated quantum adversarial examples need to undergo quality evaluation in terms of model performance and similarity before being output (Algorithm 1 line 13-15). Regarding the similarity evaluation of image data that has already been encoded into quantum states, we cannot directly employ classical methods to measure the difference or closeness between two input vectors. Serving as quantum generalizations of the classical notions, fidelity (Jozsa, 1994) and trace distance (Helstrom, 1969), which are two frequently utilized distance measures for quantum information, can assist us in efficiently calculating perturbations that fool the QNN system. As a result, we introduce the definitions of Average Fidelity Measure (AFM) and Average Trace Distance (ATD) to quantitatively assess the similarity between two quantum states, thereby evaluating the quality of the generated test samples. In fact, these metrics can also be effectively employed to assess the robustness of the model.
@@
We obtain the AFM for the QNN model $\mathcal{M}$ by calculating the sample mean of fidelity between the original inputs in $\mathcal{T}$ and their corresponding adversarial inputs. More formally:
(8) $AFM(\mathcal{M})=\frac{1}{|\mathcal{T}|}\sum_{x_{i}\in\mathcal{T}}F(|\boldsymbol{x}_{i}\rangle\langle\boldsymbol{x}_{i}|,|\boldsymbol{x}_{i}\rangle_{adv}\langle\boldsymbol{x}_{i}|_{adv}),$
@@
Figure 5 shows the visualization of fidelity between two quantum states based on POVM measurements. The minimum overlap area of the probability distributions corresponding to two quantum states obtained through POVM measurements can intuitively reflect the fidelity between them. The greater the area, the higher the fidelity.
@@
Similar to AFM, we define ATD as the average trace distance between the original and adversarial quantum samples. More precisely(More explicitly), the ATD is defined as:
(10) $ATD(\mathcal{M})=\frac{1}{|\mathcal{T}|}\sum_{x_{i}\in\mathcal{T}}D(|\boldsymbol{x}_{i}\rangle\langle\boldsymbol{x}_{i}|,|\boldsymbol{x}_{i}\rangle_{adv}\langle\boldsymbol{x}_{i}|_{adv}),$
@@
Quantum trace distance can be regarded as a generalization of classical trace distance. In fact, during simulation calculations, the results of the quantum trace distance and the $L_{1}$ distance are often very close to each other.

### v2

@@ Abstract.
Quantum Neural Network (QNN) combines the Deep Learning (DL) principle with the fundamental theory of quantum mechanics to achieve machine learning tasks with quantum acceleration. Recently, QNN systems have been found to manifest robustness issues similar to classical DL systems. There is an urgent need for ways to test their correctness and security. However, QNN systems differ significantly from traditional quantum software and classical DL systems, posing critical challenges for QNN testing. These challenges include the inapplicability of traditional quantum software testing methods to QNN systems due to differences in programming paradigms and decision logic representations, the dependence of quantum test sample generation on perturbation operators, and the absence of effective information in quantum neurons. In this paper, we propose QuanTest, a quantum entanglement-guided adversarial testing framework to uncover potential erroneous behaviors in QNN systems. We design a quantum entanglement adequacy criterion to quantify the entanglement acquired by the input quantum states from the QNN system, along with two similarity metrics to measure the proximity of generated quantum adversarial examples to the original inputs. Subsequently, QuanTest formulates the problem of generating test inputs that maximize the quantum entanglement adequacy and capture incorrect behaviors of the QNN system as a joint optimization problem and solves it in a gradient-based manner to generate quantum adversarial examples. Experimental results demonstrate that QuanTest possesses the capability to capture erroneous behaviors in QNN systems (generating 67.48%-96.05% more high-quality test samples than the random noise under the same perturbation size constraints). The entanglement-guided approach proves effective in adversarial testing, generating more adversarial examples (maximum increase reached 21.32%).
@@ 1. Introduction
Figure 1. Differences between classical DL testing (upper part) and QNN testing (lower part). In QNN system testing, the QNN components cannot extract effective neuron information as in classical DNN, and the generation of adversarial examples is also different from the classical approach.
[…]
Differences between classical DL testing (upper part) and QNN testing (lower part). In QNN system testing, the QNN components cannot extract effective neuron information as in classical DNN, and the generation of adversarial examples is also different from the classical approach.
@@
    Limited useful information can be extracted from quantum neurons. Owing to the high entanglement of quantum systems, apart from the parameters of rotation gates, we cannot guarantee meaningful information about the QNN system from the output values of individual quantum neurons, nor can we mathematically provide a specific representation similar in meaning to classical neuron outputs (Nielsen and Chuang, 2010). Valuable information often concentrates on the overall output of QNN and the corresponding measurement results. This renders existing DL testing metrics ineffective for QNN systems. Consequently, a novel evaluation metric tailored to quantum systems is imperative for QNN system testing.
In this paper, we propose QuanTest, an adversarial testing framework for QNN systems guided by quantum entanglement, to address these challenges. First, considering the unique quantum properties of the QNN system, we introduce the concept of quantum entanglement adequacy (QEA) to measure the test sufficiency of a set of test inputs to the QNN system based on the degree of activation of entanglement quantum neurons in QNN by the input quantum states. Guided by quantum entanglement adequacy, QuanTest aims to generate adversarial examples that can trigger more solution spaces (Sim et al., 2019) in QNN, thereby uncovering more erroneous behaviors in the QNN system. In particular, We first extract information from the quantum system constituted by the QNN model and quantum states. Considering the wave function collapse caused by observing the quantum systems (Nielsen and Chuang, 2010), in addition to measuring the output quantum states from the QNN model to obtain predictive information, it is necessary to retain some copies for subsequent testing to calculate the QEA and similarity metrics. For similarity evaluation, we propose two metrics, AFM and ATD, to measure the proximity between the generated adversarial examples and the original inputs. Subsequently, we formulate the problem of generating test inputs that maximize quantum entanglement adequacy of a QNN system while also exposing as many erroneous behaviors as possible as a joint optimization problem. We effectively solve this problem by obtaining perturbations in a gradient-based manner. The perturbations act directly on the original inputs in the form of a unitary operator, generating quantum adversarial examples through quantum evolution. Finally, we detect quantum adversarial examples and evaluate their quality through model performance and similarity metrics.
@@
  • We introduce quantum entanglement adequacy as the first test adequacy metric for QNN systems, and experimentally demonstrated the effectiveness of entanglement guidance in quantum adversarial testing.
@@ 2.1.1. Qubits
where $\alpha$ and $\beta$ are complex numbers, representing probability amplitudes and satisfying the condition $|\alpha|^{2}+|\beta|^{2}=1$. $\{|0\rangle,|1\rangle\}$ consists of two orthogonal 2-dimensional state vectors and is known as the computational basis in this Hilbert space. A composite system of $n$ qubits can be described by a unit complex vector in a $2^{n}$-dimensional Hilbert space. Note that the quantum state space of the composite system is the tensor product of its component system’s state spaces.
@@ 2.1.2. Quantum Logic Gates
where $U^{\dagger}$ is the conjugate transpose of $U$, and $I$ represents the identity matrix. This implies that all quantum gates must be reversible, and the dimensions of input and output must remain consistent to preserve the unitary transformation. A quantum gate acting on $n$ qubits $|\psi\rangle$ can be described by a $2^{n}\times 2^{n}$ unitary matrix $U$. The essence of this operation is to perform matrix operations on the state vector, i.e., $U|\psi\rangle$. Common quantum logic gates can be categorized based on their properties and functions into fixed gates, rotation gates, and controlled gates, among others.
@@ 2.2. Quantum Neural Network
Quantum neural networks (QNNs) are a type of neural network model based on quantum computing, with the core being parameterized quantum circuits (PQCs) (Benedetti et al., 2019) driven by variational quantum algorithms (Cerezo et al., 2021). Similar to classical feedforward neural networks, QNNs learn by updating the trainable parameters of rotation gates within the circuits through gradients——this aspect is executed by classical optimizers, and the processing and computation of the target quantum state are carried out by quantum computers. An example (Hur et al., 2022) of using a QNN model $\mathcal{M}(|\boldsymbol{x}\rangle;U_{\boldsymbol{\Theta}})$ for machine learning tasks is shown in Figure 2. The training data $\boldsymbol{x}$ is first encoded as quantum states $|\boldsymbol{x}\rangle$. Subsequently, these states are input into the QNN for computation. The trainable parameters $\boldsymbol{\Theta}$ within the PQC $U_{\boldsymbol{\Theta}}$ are updated using classical optimizers until the results meet the expected termination criteria. In this process, the PQC $U_{\boldsymbol{\Theta}}$ can be regarded as a unitary operator, and its action on the quantum state $|\boldsymbol{x}\rangle$ can be represented as $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$. The final output of the QNN can only be obtained after measuring $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$. Note that the data learned by QNN can be either classical data or quantum data, with classical data ends up encoded as quantum states (Biamonte et al., 2017). The process of quantum state encoding can be viewed as a mapping from classical dimensions to quantum dimensions, so the direct input of QNNs is always quantum states
@@ 3.1. Quantum Entanglement Adequacy Metrics
One core reason quantum computation can demonstrate potential quantum advantage in many problems is that it can construct an exponentially large quantum state space through controllable entanglement and interference (Havlíček et al., 2019). By utilizing low-depth circuits to generate quantum states with high entanglement characteristics, QNN models may produce atypical patterns that classical models cannot effectively produce. These patterns can efficiently represent solution space for tasks such as prediction and classification, capturing nontrivial correlations in the data (Schuld et al., 2020). Quantum machine learning models such as quantum circuit learning (QCL) (Mitarai et al., 2018), parameterized Hamiltonian learning (PHL) (Shi et al., 2022), and quantum kernel methods (Peters et al., 2021) achieve strong entanglement in quantum circuits for enhancing the entanglement capability of QNNs by incorporating various configurations of multi-qubit gates such as CNOT, iSWAP, and their extended variants. We can perceive that, as the quantum state used as a test input undergoes computation within a QNN, the more fully it activates the entangled quantum neurons, the higher the entanglement degree it obtained from the QNN. This leads to a more comprehensive coverage of the solution space in the QNN. Therefore, we propose a method to calculate the QEA of QNNs. This is achieved by measuring the entanglement of both the test input quantum state and the resulting output quantum state after undergoing unitary operations within the QNN. The difference in measurement outcomes is used to represent the QEA of that particular test case.
@@
Given a fully trained QNN model $\mathcal{M}$ and a test input set $\mathcal{T}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$, the action of $\mathcal{M}$ on the input quantum state $|\boldsymbol{x}\rangle$ can be equivalently represented as the multiplication of $|\boldsymbol{x}\rangle$ and a unitary matrix $U_{\boldsymbol{\Theta}}$, where $|\boldsymbol{x}\rangle$ represents the quantum state that encodes the information of $\boldsymbol{x}\in\mathcal{T}$, and $\boldsymbol{\Theta}$ represents the learned and fixed gate parameters in $\mathcal{M}$. Note that the result output of $U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle$ is a quantum state with the same dimension as $|\boldsymbol{x}\rangle$. Then, the quantum entanglement adequacy (QEA) can be expressed as:
@@
Figure 4. Example of QEA computation for two quantum states in the QNN. The nodes in gray denote quantum entanglement layers that are not sufficiently activated, while aqua nodes represent those that are adequately activated.
[…]
Example of QEA computation for two quantum states in the QNN. The nodes in gray denote quantum entanglement layers that are not sufficiently activated, while aqua nodes represent those that are adequately activated.
@@
3.2. QuanTest Algorithm
3.2.1. Joint Optimization
It consists of two parts, where the first part $k\cdot Q(U_{\boldsymbol{\Theta}}|\boldsymbol{x}\rangle)-Q(|\boldsymbol{x}\rangle)$ is to generate test inputs that maximize QEA, and the second part $f(\mathcal{M}(|\boldsymbol{x}\rangle;U_{\boldsymbol{\Theta}});\mathcal{A})$ is to generate test inputs that trigger as many incorrect behaviors in the QNN as possible. The hyperparameter $\lambda$ is used to balance the two optimization objectives during the joint optimization process.
@@ 3.2.1. Joint Optimization
10:         obj = [];
11:         obj.append(obj_incorrect($|\boldsymbol{x}\rangle_{out}$, strategy));
12:         obj.append(w $\cdot$ obj_entanglement($|\boldsymbol{x}\rangle_{adv}$, $|\boldsymbol{x}\rangle_{out}$, k));
@@
We treat these two parts as a joint optimization problem and maximize this objective function (Algorithm 1 line 8-11).
@@
  • Maximizing incorrect behaviors. Modular construction enables QuanTest to conveniently select different quantum-adapted adversarial attack strategies $\mathcal{A}$. The determination of this strategy dictates the prediction information to be further calculated after measuring the output quantum state during the information extraction. For instance, the quantum-adapted FGSM strategy requires providing loss, while the quantum-adapted DLFuzz strategy needs category probabilities as input. In this part, QuanTest aims to modify the input in the form of adversarial attacks, leading the model to exhibit incorrect behavior.
All terms of the jointly optimized objective are differentiable. QuanTest can iteratively modify the input $|\boldsymbol{x}\rangle$ by using gradient ascent to maximize these terms. The computed gradient $\frac{\partial\text{obj}}{\partial|\boldsymbol{x}\rangle}$ will become the perturbation, which is subsequently transformed into a perturbation operator on a quantum circuit for generating adversarial examples.
@@ 3.2.3. Quality Evaluation
The generated quantum adversarial examples need to undergo quality evaluation in terms of model performance and similarity before being output as test samples (Algorithm 1 line 13-15). A qualified test sample should first be able to induce erroneous behavior in the model, such as causing a change in the predicted labels of the classification model. Regarding the similarity evaluation of image data that has already been encoded into quantum states, we cannot directly employ classical methods to measure the difference or closeness between two input vectors. Serving as quantum generalizations of the classical notions, fidelity (Oh et al., 2019) and trace distance (Helstrom, 1969) are two frequently utilized distance measures for quantum information. They have also been used in some quantum machine learning work (Lu et al., 2020; Liao et al., 2021; Anil et al., 2024) to evaluate the similarity between quantum states. Here, we choose to use these measures to calculate perturbations that fool the QNN system. As a result, we introduce the definitions of Average Fidelity Measure (AFM) and Average Trace Distance (ATD) to quantitatively assess the similarity between the original input quantum state and the quantum test samples, thereby evaluating the quality of the generated test samples. In fact, these metrics can also be effectively employed to assess the robustness of the model.
@@
We obtain the AFM by calculating the sample mean of fidelity between the original inputs in $\mathcal{T}$ and their generated test samples. More formally:
(8) $AFM(\mathcal{T})=\frac{1}{|\mathcal{T}|}\sum_{x_{i}\in\mathcal{T}}F(|\boldsymbol{x}_{i}\rangle\langle\boldsymbol{x}_{i}|,|\boldsymbol{x}_{i}\rangle_{adv}\langle\boldsymbol{x}_{i}|_{adv}),$
@@
The minimum overlap area of the probability distributions corresponding to two quantum states obtained by an optimal positive-operator valued measure (POVM) can intuitively reflect the fidelity between them (Oh et al., 2019). Figure 5 shows the visualization of fidelity between two quantum states based on POVM measurements. The greater the area, the higher the fidelity.
@@
Similar to AFM, we define ATD as the average trace distance between the original inputs and their generated test samples. More precisely, the ATD is defined as:
(10) $ATD(\mathcal{T})=\frac{1}{|\mathcal{T}|}\sum_{x_{i}\in\mathcal{T}}D(|\boldsymbol{x}_{i}\rangle\langle\boldsymbol{x}_{i}|,|\boldsymbol{x}_{i}\rangle_{adv}\langle\boldsymbol{x}_{i}|_{adv}),$
@@
Quantum trace distance can be regarded as a generalization of classical trace distance. It represents an achievable upper bound on the total variation distance between probability distributions arising from measuring two quantum states (Liao et al., 2021; Nielsen and Chuang, 2010). In fact, during simulation calculations, the results of the quantum trace distance and the $L_{1}$ distance are often very close to each other.